A car traveling at 20.0 m/s starts to decelerate at a constant rate. It comes to a stop in 7.5 seconds. What is the cars acceleration


1
Expert's answer
2021-02-18T18:39:16-0500
"a=\dfrac{v-v_0}{t}=\dfrac{0-20\ \dfrac{m}{s}}{7.5\ s}=-2.67\ \dfrac{m}{s^2}."

The sign minus means that the car decelerates.